There exists a field with..... elements


1
Expert's answer
2021-02-04T07:59:05-0500

There exists a field with two elements. An example of such a field is "(\mathbb Z_2,+,\cdot )" where "\mathbb Z_2=\{0,1\}" and "0+0=1+1=0, \ 0+1=1+0=1,\ 0\cdot 0 = 0\cdot 1 = 1\cdot 0 =0,\ 1\cdot 1=1."